OpenStudy (anonymous):

Lily purchases a ring for $990 that appreciates at a rate of 7% per year. What will the value of the ring be in 10 years?

OpenStudy (tkhunny):

This is a single payment annual accumulation. You need to get VERY GOOD at these. Everything else is harder! \(990(1.07)^{10}\)

OpenStudy (anonymous):

Alright, thank you. I tend to struggle with turning the percents into multiplied decimals. How does appreciating at 7% mean multiply by 1.07? (I understand the exponent part)

OpenStudy (tkhunny):

7% = 0.07 = 7/100 1 + 7% = 1.07 = 107/100